OpenStudy (anonymous):

Quick Geometric Series Question. I dont understand how to get r in the equation for finding the sum. sum = a/(1-r) I understand how to get a though.

mathslover (mathslover):

sum * (1-r) = a 1-r = a/(sum) r = 1- ( a/ sum )

OpenStudy (anonymous):

I mean without knowing the sum. For example, ive attached a problem. a = 10/11

OpenStudy (anonymous):

r is your common ratio.

OpenStudy (anonymous):

i know i just dont understand how to get that. I can usually guess it but i would like to know how to find it.

OpenStudy (anonymous):

You know, if there is a way to find r without knowing the sum, I do not know it. I hope someone else does though because now I'm curious as well. I'm sorry, I wish I could help!

OpenStudy (anonymous):

if some one could just help me get r in this problem that would be fine.